OpenStudy (anonymous):

Kenny goes shopping for school clothes during a tax−free weekend, and he saves 6% sales tax. Use a percent proportion to determine how much he would save if he bought clothes worth $300. Which of the following is true? The solution is $18, and is a denominator in the proportion. The solution is $18, and is a numerator in the proportion. The solution is $5,000 , and is a denominator in the proportion. The solution is $5,000 , and is a numerator in the proportion.

OpenStudy (anonymous):

@ankit042 you know?

OpenStudy (ankit042):

find 6% of 300 0.06*300 =18% x/300 =6/100